OpenStudy (anonymous):

Given the equation -4(2-y)+5x=6(x-y). In order to graph this linear equation using a graphing calculator, it must be solved for y. Solve for y.

OpenStudy (anonymous):

use the Distributive Law, - 8 + 4y + 5x = 6x - 6y Add 6y to both sides, we get - 8 + 10y + 5x = 6x Subtract 5x from both sides, we get -8 + 10y = x Add 8 to both sides, we get 10y = x + 8 Finally, divide both sides by 10, we get y = (x + 8)/10 Make sure that you put the x+8 in parenthesis or you will be wrong with a different graph!